The director of social services of a county has learned that the state has mandated additional information requirements. This will place an additional burden on the agency. The director has identified three acceptable alternatives to handle the increased workload. One alternative is to reassign present staff members, the second is to hire and train two new workers, and the third is to redesign current practice so that workers can readily collect the information with little additional effort. An unknown factor is the caseload for the coming year when the new data will be collected on a trial basis. The estimated costs for various options and caseloads are shown in the following table:

CASELOAD
Moderate High Very High
Reassign staff $ 50 * 70 60
New staff 70 70 70
Redesign collection 85 90 95

*Cost in $ thousands.


After contemplating the caseload question, the director of social services has decided that reasonable caseload probabilities are 0.20 for moderate, 0.30 for high, and 0.50 for very high.

a. Which alternative will yield the minimum expected cost?

  • New Staff

  • Reassign

  • Redesign



c. Determine the expected value of perfect information using an opportunity loss table. (Round your answers to the nearest whole number. Leave no cells blank - be certain to enter "0" wherever required.)

Opportunity loss table
Moderate($) High($) Very High($) EOL($)
Reassign
New Staff
Redesign
(0.20) (0.30) (0.50)


Expected value of perfect information           $
Answer #1

Using treeplan plugin for excel ,we draw the decision tree as below.

a.The alternative with minimum expected cost is Reassign staff and the expected cost is $ 61 thousands

b.

Expected value of perfect information =Expected value with perfect information-Expected value without perfect information

=0.2 *50 +0.3 *70+0.5*60 -EVwoPI

=10+21+30-EVwoPI

=61-EVwoPI

=61-EVwoPI

But from the decision tree,we have EVwoPI=61

So EV of perfect information=61-61=0

EV of perfect information=0